### v1

# Game-Theoretic Coordination For Time-Critical Missions of UAV Systems

Abstract
Cooperative missions involving Unmanned Aerial Vehicles (UAVs) in dynamic environments pose significant challenges in ensuring both coordination and agility. In this paper, we introduce a novel game-theoretic approach for time-critical missions, where each UAV optimizes a cost function that incorporates temporal and mission-specific constraints. The optimization is performed within a one-dimensional domain, significantly reducing the computational cost and enabling real-time application to complex and dynamic scenarios. The framework is distributed in structure, allowing to achieve global, system-wide coordination (a Nash equilibrium) by using only local information. For ideal systems, we prove the existence and exponential stability of the Nash equilibrium. Furthermore, we invoke model predictive control (MPC) for non-ideal scenarios. In particular, we propose a discrete-time optimization approach that tackles path-following errors and communication failures, ensuring reliable and agile performance in dynamic and uncertain environments. Simulation results demonstrate the effectiveness and agility of the approach in ensuring successful mission execution across diverse scenarios. Experiments using a motion capture system provide further validation under realistic conditions.
I Introduction
Recent advancements in technology have significantly expanded the capabilities of Unmanned Aerial Vehicle (UAV) systems, enabling them to perform a wide range of tasks [27, 35]. Particularly, cooperative UAV systems have been extensively deployed across diverse domains, including civilian applications such as surveillance, environmental monitoring, and air traffic management, as well as military operations like swarm-based attack-defense scenarios. Notable examples include cooperative forest fire monitoring and suppression [28, 34], multiple moving targets surveillance [14], to name a few. Despite their broad range of applications, cooperative UAV systems face operational limitations and challenges like battery endurance, payload carrying capability, flight autonomy, path planning, path following, and achieving reliable cooperation.
Figure 1: Snapshot of the flight experiment demonstrating four UAVs coordinating and avoiding collisions in the presence of non-ideal communication. The added trail effect highlights the flight path of UAVs, a circular symbol representing the balance between two opposing forces, yin and yang.
Effective cooperation among a system of UAVs is crucial for accomplishing complex tasks that exceed the capabilities of a single UAV. Applications such as simultaneous target tracking [3], formation flying [23], and large-scale area mapping and monitoring [32] heavily rely on the coordinated efforts of multiple UAVs. Achieving reliable cooperation, however, is inherently challenging due to factors such as communication delays, limited computational resources, variations in UAV dynamics, disturbances, and unpredictable events. These challenges highlight the need for efficient and adaptive strategies to ensure system-wide collaboration in dynamic and uncertain environments. This paper addresses the challenge of time-critical cooperation through a novel game-theoretic method.
Inspired by the pioneering work of Leslie Lamport on time synchronization [21], we formulate the time-critical coordination problem as a “time synchronization” problem (see Section II for more details). More precisely, following the framework of cooperative path-following in [2, 15, 18, 9], we decouple the space and time in the general problem formulation, which reduces the coordination problem to a one-dimensional consensus problem, thereby significantly lowering the computation costs. In [15, 17, 8, 18, 19], exponential stability of time coordination was established with a PI controller for networks connected in an integral sense. In this paper, we formulate the coordination problem as a game, and we use the virtual time (see [18]) as a consensus parameter for UAVs that engage in a game to reach an agreement on this parameter. As compared to the PI controller used in earlier work for coordination, we now provide a general approach to handle operational requirements and time-varying objectives. Furthermore, certain constraints, such as UAV speed limits or acceleration limits, may be directly incorporated into the optimization problem to ensure the feasibility of the provided solution. The optimization problem also enables the direct incorporation of complex, time-varying mission specifications into the optimization framework (e.g., loss of energy efficiency that might require mission modification or accommodating new collision avoidance requirements not accounted for in the apriori mission planning, as demonstrated in Section III-B, Remark 2). Figure 2 shows the architecture of the solution. The main difference from prior papers [2, 15, 18], which used a PI control law for coordination without considering optimality, lies in the block on Optimal Temporal Coordination, where a game theoretic formulation is considered for the UAVs to reach agreement on the consensus variables. The existence of a Nash equilibrium is shown and its exponential stability is proven.
Next, we highlight the main contributions of this paper:
  • We propose a novel game-theoretic coordination algorithm, where the Nash equilibrium of the strategic interactions among agents ensures system-wide coordination and successful task execution.
  • We prove the existence of a Nash equilibrium and establish the exponential stability under ideal conditions.
  • Through extensive simulations and flight experiments, we demonstrate the effectiveness of the algorithm for various coordination tasks in UAV systems.
In contrast to the classical Nash equilibrium existence results (see, for example, [24] and [37]), we prove the existence of continuous-time, infinite horizon Nash equilibrium in infinite-dimensional action domain with constraints. A similar result, like the existence of a Nash equilibrium with bounded control inputs and stability of the solution, was established in [36] using Lyapunov stability techniques. In this paper, we prove the existence and exponential stability of a Nash equilibrium with constraints, relying on Euler-Lagrange equations. For the Nash equilibrium analysis, Euler-Lagrange equations were used in [10] in the case of a single integrator. Here, we consider double integrator, where the corresponding system of Euler-Lagrange equations has a higher degree, and we study this system by explicitly solving it. Regarding Nash equilibrium seeking algorithms, most of the known results assume a fixed communication network (see, for example [25, 33, 13]). In contrast, leveraging MPC, we propose an algorithm that converges to Nash equilibrium even in the case of time-varying network.
The remainder of this paper is structured as follows: Section II introduces the necessary notations and provides an overview of the preliminaries, including path planning, path following and coordination. Furthermore, in Section II, we introduce an infinite horizon game problem with discount factor. In Section III, by leveraging the system of Euler-Lagrange equations (a coupled fourth-order differential equations), we prove the existence of a Nash equilibrium for the corresponding unconstrained problem under ideal conditions. Moreover, we derive an explicit expression for the Nash equilibrium and prove the exponential stability of the solution; see Proposition 2. Eventfully, examining the coefficients of the explicit solution, we prove that there exists a discount rate such that the solution to the unconstrained problem is also solution to the constrained one; see Theorem 3. Section III-B outlines the general algorithm derived from the theoretical results, and Section IV validates the algorithm’s reliability and performance through simulations in challenging and realistic scenarios. Finally, Section V describes the flight experiments conducted with small UAVs. The video footage showcasing the simulations and the flight experiments is provided in the supplementary materials.
II Preliminaries
We consider a class of time-critical cooperative missions that can be formulated as “time synchronization” problems (see, for example, [21, 18]). We assume that for a given mission a path planner generates a desired and feasible trajectory (possibly optimal for the mission) for each UAV. Next, we assume that each UAV is equipped with a robust low-level controller for tracking purposes. The coordination task is then reduced to a “time synchronization” problem.
Figure 2: Game-theoretic cooperative path-following of a system of UAVs: the block on Optimal Temporal Coordination uses a game-theoretic formulation and ensures existence and convergence of Nash equilibrium.
II-A Path Planning and Following
Consider a system of $N$ UAVs involved in the mission, $N\in\mathbb{N}$. For each UAV the desired trajectory is given as a function $x_{d,i}:[0,t_{d,i}^{*}]\rightarrow\mathbb{R}^{3}$, $t_{d,i}^{*}>0$. For all $t\in[0,t_{d,i}^{*}]$, the trajectories satisfy the constraints defined by the $i^{\rm th}$ UAV’s dynamics:
$0\leqslant v^{i}_{\min}<v^{i}_{d,\min}\leqslant\left|\left|\dot{x}_{d,i}(t)\right|\right|\leqslant v^{i}_{d,\max}<v^{i}_{\max},$ (1)
where $v^{i}_{\min}$ and $v^{i}_{\max}$ are the $i^{\rm th}$ vehicle’s possible minimum and maximum speeds, respectively, while, $v^{i}_{d,\min}$ and $v^{i}_{d,\max}$ are the vehicle’s minimum and maximum speeds for the given mission. Similarly,
$\left|\left|\ddot{x}_{d,i}(t)\right|\right|\leqslant a^{i}_{d,\max}<a^{i}_{\max},$ (2)
with $a^{i}_{\max}$ being the $i^{\rm th}$ UAV’s possible maximum acceleration and $a^{i}_{d,\max}$ being the maximum acceleration for the mission.
Time-critical mission specifications are encoded into the desired trajectories of the UAVs by the planner. By properly defining these trajectories, it is possible to design missions with varying characteristics, such as simultaneous arrival and/or sequential autolanding. The analysis of such systems are similar, and hence, here, we consider only the simultaneous arrival case. A simultaneous arrival mission is defined as:
$t_{d,i}^{*}=t_{d}^{*},\quad i=1,\dots,N.$
The trajectories must maintain spatial separation throughout the mission:
$\min_{\begin{subarray}{c}i,j=1,\ldots,n\\ i\neq j\end{subarray}}\left\|x_{d,i}(t)-x_{d,j}(t)\right\|^{2}\geqslant E^{2}>0,\quad\text{for all }t\geqslant 0.$ (3)
Remark 1. In Section III-B, we introduce an additional collision avoidance term, which allows to remove the spatial separation constraints in (3).
Various approaches exist for trajectory generation, each for specific objectives; e.g., optimal control-based methods (minimizing energy consumption or path length), waypoint-based trajectory generation (connects waypoints using smooth polynomials), and minimum-snap trajectory generation algorithm (see [5, 6, 11, 26]).
We assume that each UAV is equipped with a path-following controller that ensures that the UAV follows its desired trajectory, $x_{d,i}(t)$, or the updated trajectory based on coordination. Examples include PID controllers [31], geometric controllers [22], and adaptive path-follower [38], to name a few.
II-B Time Coordination for Simultaneous Arrival
To formulate the problem of simultaneous arrival of UAVs, we follow [8, 18]. Towards that end, let $\gamma_{i}:\mathbb{R}^{+}\rightarrow[0,t_{d}^{*}]$, $i=1,\dots,N$, map the actual (clock) time $t$ to the planned mission time $t_{d}$, that is, $\gamma_{i}(t)=t_{d}$. In [8, 18], $\gamma_{i}$ is referred to as virtual time, which allows to redefine the desired trajectory of the $i^{\rm th}$ UAV as:
$x_{\gamma,i}\left(t\right)=x_{d,i}\left(\gamma_{i}(t)\right).$ (4)
Note that by this reparametrization of the desired trajectories, we gain control over the speed of each UAV, and hence, its mission execution time. Specifically,
  • When $\dot{\gamma}_{i}(t)=1$, the UAV travels at the desired pace.
  • When $\dot{\gamma}_{i}(t)>1$, the UAV moves faster than the desired pace.
  • When $\dot{\gamma}_{i}(t)<1$, the UAV moves slower than the desired pace.
Thus, $\gamma_{i}$ serves the role of a consensus parameter in our problem. Next, based on the physical limitations and mission requirements of the UAVs, as defined in (1) and (2), we derive general bounds for the consensus parameter of each UAV. Specifically, from (4), we have
$\begin{split}\dot{x}_{\gamma,i}(t)&=\dot{x}_{d,i}\left(\gamma_{i}(t)\right)\dot{\gamma}_{i}(t),\\ \ddot{x}_{\gamma,i}(t)&=\ddot{x}_{d,i}\left(\gamma_{i}(t)\right)\dot{\gamma}_{i}(t)^{2}+\dot{x}_{d,i}\left(\gamma_{i}(t)\right)\ddot{\gamma}_{i}(t).\end{split}$ (5)
The first equation along with (1) implies that the minimum and the maximum values of all admissible parameters $\gamma_{i}$ must satisfy the following constraint:
$\frac{v^{i}_{\min}}{v^{i}_{d,\min}}\leqslant\dot{\gamma}^{i}_{\min}\leqslant\dot{\gamma}_{i}(t)\leqslant\dot{\gamma}^{i}_{\max}\leqslant\frac{v^{i}_{\max}}{v^{i}_{d,\max}}.$ (6)
These inequalities ensure that the UAVs maintain forward motion throughout the mission (due to $\dot{\gamma}_{i}(t)\geqslant 0$ condition). On the other hand, by (2) and the second equation of (5) it follows that $\dot{\gamma}^{i}_{\max}=\max_{t}\{\dot{\gamma}_{i}(t)\}$ and $\ddot{\gamma}^{i}_{\max}=\max_{t}\{|\ddot{\gamma}_{i}(t)|\}$ should satisfy
$\ddot{\gamma}^{i}_{\max}v^{i}_{d,\max}+(\dot{\gamma}^{i}_{\max})^{2}a^{i}_{d,\max}\leqslant a^{i}_{\max}.$ (7)
Note that since $v^{i}_{d,\max}<v^{i}_{\max}$ and $a^{i}_{d,\max}<a^{i}_{\max}$, there exist $\gamma_{i}$ parameters such that inequalities in (6) and (7) are satisfied.
The time coordination of the system is achieved when the consensus parameters are synchronized (see [18] and references within); that is,
$\gamma_{i}(t)-\gamma_{j}(t)=0,\quad\text{for all }i,j\in\{1,2,\dots,N\}.$ (8)
This condition ensures that all UAVs reach their respective goal positions simultaneously. Furthermore, to ensure that the UAVs maintain a predefined desired speed profile, the derivatives of the virtual times should satisfy:
$\dot{\gamma}_{i}(t)-1=0,\quad\text{for all }i\in\{1,2,\dots,N\}.$ (9)
II-C Problem Formulation
Next, we introduce a game-theoretic approach that facilitates time-critical cooperative mission execution. For the theoretical analysis, here, we consider the path-following kinematics (ideal path-following) and a completely connected communication network. While these assumptions simplify the theoretical analysis (see Section III-A), they do not limit the practical applicability of the proposed approach. In Section III-B, we introduce an algorithm that incorporates path-following errors and communication failures, enabling application in real-world scenarios. As shown in [7, 16], by using cascaded inner-loop outer-loop structure for flight control applications, the uncertainties in system dynamics can be handled by a variety of robust control methods.
We begin by introducing the admissible set of the consensus parameter (virtual time) for each UAV. We have $\gamma_{i}(0)=\gamma^{0}_{i}\geqslant 0$, for $i=1,\dots,N$. Along with the inequalities in (6), (7), we define the admissible sets:
$\mathcal{A}^{0}_{i}:=\Big{\{}\gamma_{i}\in H^{1}_{w,\alpha}((0,\infty)):\gamma_{i}(0)=\gamma^{0}_{i}\geqslant 0,\,\dot{\gamma}_{i}(0)=1\Big{\}},$ (10)
and
$\mathcal{A}^{2,\alpha}_{i}:=\Big{\{}\gamma_{i}\in\mathcal{A}^{0}_{i}:\dot{\gamma}_{i}\geqslant 0,\,\|\dot{\gamma}_{i}\|_{L^{\infty}}\leqslant V^{i}_{1},\,\|\ddot{\gamma}_{i}\|_{L^{\infty}}\leqslant V^{i}_{2}\Big{\}},$
where $H^{1}_{w,\alpha}((0,\infty))$ is weighted Sobolev space (for more details see Appendix VIII) with weight $e^{-\alpha t}$
$\begin{split}H^{1}_{w,\alpha}((0,\infty)):=\Bigg{\{}g\in H^{1}_{loc}((0,\infty)):\\ \int_{0}^{\infty}e^{-\alpha t}(g^{2}+\dot{g}^{2}){\rm d}t<\infty\Bigg{\}}.\end{split}$ (11)
Problem 1
Consider a system of $N$ UAVs. Each UAV (agent) over $\gamma_{i}\in{\mathcal{A}}^{2,\alpha}_{i}$ seeks to minimize
$\begin{split}J_{i}(\gamma_{i},\gamma_{-i})=\int_{0}^{\infty}e^{-\alpha t}\Bigg{(}\left(\dot{\gamma}_{i}-1\right)^{2}+\sum_{j=1}^{N}(\gamma_{i}-\gamma_{j})^{2}+\ddot{\gamma}_{i}^{2}\Bigg{)}{\rm d}t,\end{split}$ (12)
where $\gamma_{-i}=(\gamma_{1},\dots,\gamma_{i-1},\gamma_{i+1},\dots,\gamma_{N})$.
The cost function in (12) corresponds to simultaneous arrival and maintaining a desired mission pace with the following terms:
  • A pace penalty: $(\dot{\gamma}_{i}-1)^{2}$, penalizes deviation from the desired pace.
  • A coordination penalty: $(\gamma_{i}-\gamma_{j})^{2}$ penalizes for the discrepancies with neighbors.
  • A control effort penalty: $(\ddot{\gamma}_{i})^{2}$ penalizes excessive control inputs.
We refer to the solution of Problem 1 as Nash equilibrium.
Definition 1
The vector function $\gamma^{*}=(\gamma^{*}_{1},\dots,\gamma^{*}_{N})$ is Nash equilibrium of Problem 1, if for all $\gamma_{i}\in{\mathcal{A}}^{2,\alpha}_{i}$
$J_{i}(\gamma^{*}_{i},\gamma^{*}_{-i})\leqslant J_{i}(\gamma_{i},\gamma^{*}_{-i}).$ (13)
III Main Approach
In this section, we prove the existence and exponential stability of the solution to Problem 1. Furthermore, we introduce a MPC based algorithm that approximates the solution of Problem 1.
III-A Existence of Solution and Exponential Stability
Under ideal system conditions, we isolate and analyze the core features of the game-theoretic approach. More precisely, to establish the existence and exponential stability of the solution to Problem 1, we first prove the existence and exponential stability of the solution to the corresponding unconstrained problem. Then, we determine the parameter $\alpha$ (discount rate) such that the solution to the unconstrained problem is also a solution to the constrained problem. We then conclude that, for the chosen $\alpha$, Problem 1 has a solution and that the solution is exponentially stable.
The unconstrained problem can be formulated as follows:
Problem 2
Consider a system of $N$ UAVs. Each UAV (agent) seeks to minimize
$\begin{split}\int_{0}^{\infty}e^{-\alpha t}\Bigg{(}\dot{\gamma}_{i}^{2}+\sum_{j=1}^{N}(\gamma_{i}-\gamma_{j})^{2}+\ddot{\gamma}_{i}^{2}\Bigg{)}{\rm d}t\end{split}$ (14)
over $\gamma_{i}\in\mathcal{B}^{0}_{i}=\{\gamma_{i}\in H^{1}_{w,\alpha}((0,\infty)):(\gamma_{i}+t)\in\mathcal{A}^{0}_{i}\}$.
The following result proves the existence and exponential stability of the solution to Problem 2 and provides an explicit expression.
Proposition 2
Let $\alpha>0$. Then, there exists $\bm{\gamma}^{*}=(\gamma_{1}^{*},\dots,\gamma_{N}^{*})\in\prod_{j=1}^{N}\mathcal{B}^{0}_{j}$ solving Problem 2. Moreover, the solution has the following explicit form
$\gamma_{i}^{*}(t)=H^{1}_{i}+H^{3}_{i}e^{\mu_{3}t}+e^{\mu_{1}t}\Big{(}C_{i}^{1}\cos(\nu_{1}t)+C_{i}^{2}\sin(\nu_{1}t)\Big{)},$ (15)
where $\mu_{1},\mu_{3}<0$, and the constants $H^{1}_{i},H_{i}^{3},C_{i}^{1},\nu_{1},C_{i}^{2}$ only depend on $\alpha$ and initial conditions.
Proof:
Here, we present the general idea of the proof; for more details see Appendix IX.
Suppose that $\bm{\gamma}^{*}=(\gamma_{1}^{*},\dots,\gamma_{N}^{*})\in\prod_{j=1}^{N}\mathcal{B}^{0}_{j}$ is a solution to Problem 2. Then, by the definition of Nash equilibrium and the convexity of the integrand of (14) it follows that $\gamma_{i}^{*}$ is the unique minimizer of the following optimization problem
[…]
To examine the behavior of $\gamma_{i}^{*}$, $i=1,\dots,N$, at infinity, we consider Euler-Lagrange equations of (16)
$\gamma_{i}^{(4)*}-2\alpha\dddot{\gamma}_{i}^{*}+(\alpha^{2}-1)\ddot{\gamma}_{i}^{*}+\alpha\dot{\gamma}_{i}^{*}+\sum_{j=1}^{N}(\gamma_{i}^{*}-\gamma_{j}^{*})=0,$ (17)
with the transversality conditions
$\begin{split}&\lim_{T\to\infty}e^{-\alpha T}\ddot{\gamma}_{i}^{*}(T)=0,\\ &\lim_{T\to\infty}e^{-\alpha T}\Big{(}\dot{\gamma_{i}}^{*}(T)+\alpha\ddot{\gamma}_{i}^{*}(T)-\dddot{\gamma}_{i}^{*}(T)\Big{)}=0.\end{split}$ (18)
Using (18) and the boundary conditions (10), we solve the system of Euler-Lagrange equations in (17) and obtain (15).
To prove the existence of the solution to Problem 2, we repeat the steps above in the opposite direction. Specifically, since $\gamma_{i}^{*}$ in (15) is solving the Euler-Lagrange equation in (17), and the variational problem in (16) is quadratic, we have that $\gamma_{i}^{*}$ is the minimizer of (16). Subsequently, $\bm{\gamma}^{*}$ is a Nash equilibrium of Problem 2. ∎
Next, we prove the existence and exponential stability of the solution to the constrained problem, i.e. Problem 1.
Theorem 3
For any initial condition, $\gamma^{0}=(\gamma_{1}^{0},\dots,\gamma_{N}^{0})$, and physical constraints on the UAVs ($V^{i}_{1}$ and $V_{2}^{i}$), there exists $\alpha>0$ such that Problem 1 has a solution, and it is exponentially stable.
Proof:
In the detailed proof of Proposition 2 (see Appendix IX), we keep track of the exact dependence of the constants $H^{1}_{i},H_{i}^{3},\mu_{1},\mu_{3},\nu_{1},C_{i}^{1},C^{2}_{i}$ on the parameter $\alpha$. Using these dependencies, we have
$\begin{split}H^{1}_{i}=\gamma_{i}^{0}+O(\tfrac{1}{\alpha^{2}}),\quad H_{i}^{3}=O(\tfrac{1}{\alpha^{3}}),\quad C^{1}_{i}=O(\tfrac{1}{\alpha^{2}}),\\ C^{2}_{i}=O(\tfrac{1}{\alpha}),\quad\mu_{3}=O(\tfrac{1}{\alpha}),\quad\mu_{1}=O(\tfrac{1}{\alpha}),\quad\nu_{1}=O(\tfrac{1}{\alpha}),\end{split}$ (19)
as $\alpha\to\infty$. On the other hand, the explicit solution to Problem 2 (see (15)) is a combination of uniformly bounded functions in $t$
$\cos(\nu_{1}t),\quad\sin(\nu_{1}t),\quad e^{\mu_{1}t},\quad e^{\mu_{3}t}.$
This with (19) implies that taking large enough $\alpha$ for the explicit solution to Problem 2, we obtain $(\gamma_{i}^{*}-t)\in\mathcal{A}^{2,\alpha}_{i}$. Therefore, $(\bm{\gamma}^{*}-t)$ is a solution to the constrained problem, Problem 1, as well. By Proposition 2, it follows that the solution is exponentially stable. ∎
III-B Algorithm
In this part, we develop an algorithm using Problem 1 that leads to time coordination among UAVs. The algorithm uses MPC, resulting in a finite-dimensional optimization problem, which serves as an approximation to the solution to Problem 1, while accommodating real-world challenges such as communication constraints, path-following errors and dynamic environments.
In the formulation of Problem 1, the communication network is static and completely connected; i.e. each UAV is connected to all others involved in the mission. The proposed algorithm relaxes these assumptions. We assume that the communication network remains static during each MPC step but may change between steps. The communication network change is incorporated into the cost function. Furthermore, the algorithm considers the path-following errors in the constraints that may arise when each UAV follows its desired trajectory operating in an uncertain and dynamic environment.
Problem 3
Let $h>0$ be the time step. To approximate the virtual time at time $t_{k}=kh$, each UAV solves the following minimization problem:
$\begin{cases}\min_{y^{k}_{i}}J_{i}\left(y^{k}_{i},\bar{s}^{k}_{-i}\right)\\ s^{k}_{i\tau+1}=s^{k}_{i\tau}+h\ell^{k}_{i\tau}+\dfrac{h^{2}}{2}u^{k}_{i\tau},\quad\tau=0,\dots,K-1,\\ \ell^{k}_{i\tau+1}=\ell^{k}_{i\tau}+hu^{k}_{i\tau},\quad\tau=0,\dots,K-1,\\ s^{k}_{i0}=s^{k-1}_{i1}-\alpha_{i}^{k},\quad\ell^{k}_{i0}=\ell^{k-1}_{i1},\\ \dot{\gamma}^{i}_{\min}\leqslant\ell^{k}_{i\tau}\leqslant\dot{\gamma}^{i}_{\max},\quad\tau=0,\dots,K,\\ \left|u^{k}_{i\tau}\right|\leqslant\ddot{\gamma}^{i}_{\max},\quad\tau=0,\dots,K-1,\end{cases}$
for $k=1,2,\dots$. The optimization variables are
$y^{k}_{i}=[s^{k}_{i},\ell^{k}_{i},u^{k}_{i}],$
where
$\begin{split}s^{k}_{i}=[s^{k}_{i0},s^{k}_{i1},\dots,s^{k}_{iK}],\\ \ell^{k}_{i}=[\ell^{k}_{i0},\ell^{k}_{i1},\dots,\ell^{k}_{iK}],\\ u^{k}_{i}=[u^{k}_{i0},u^{k}_{i1},\dots,u^{k}_{iK-1}].\end{split}$
The cost function $J_{i}$ is defined as follows:
$J_{i}\left(y^{k}_{i},\bar{s}^{k}_{-i}\right)=\sum_{\tau=1}^{K}\left(\ell^{k}_{i\tau}-1\right)^{2}+F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i}\right)+\sum_{\tau=0}^{K-1}{u^{k}_{i\tau}}^{2},$ (20)
with
$\bar{s}^{k}_{-i}=\left[\bar{s}^{k}_{1},\dots,\bar{s}^{k}_{i-1},\bar{s}^{k}_{i+1},\dots,\bar{s}^{k}_{N}\right]$ (21)
and
$\begin{split}F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i}\right)=\sum_{j\in\mathcal{N}_{ik}}\sum_{\tau=1}^{K}\left(s^{k}_{i\tau}-\bar{s}^{k}_{j\tau}\right)^{2},\end{split}$ (22)
where $\mathcal{N}_{ik}$ denotes the neighborhood of $i^{\rm th}$ UAV, which is the set of UAVs that communicate with $i^{\rm th}$ one at time $t_{k}$.
All terms of the cost function in (20) have been discussed in the continuous case; see Section II-C. The newly added correction term $\alpha_{i}^{k}$ appears in initial conditions and takes into account the path-following error:
$\alpha_{i}^{k}=\alpha_{i}^{k}(x_{i}(t_{k}))=\beta\dfrac{e_{x,i}(t_{k})^{T}\dot{x}_{\gamma,i}(t_{k})}{\left|\left|\dot{x}_{\gamma,i}(t_{k})\right|\right|+\delta},$ (23)
where $\beta$ and $\delta$ are positive parameters. The correction term $\alpha_{i}^{k}$ is positive if the UAV’s actual location projection on the desired trajectory is ahead of the desired position $x_{\gamma,i}(t_{k})$ and negative otherwise. Including $\alpha_{i}^{k}$ in initial conditions causes the UAV to slow down to reduce the forward overshoot and to accelerate to catch up with the desired trajectory.
At next time instance $t_{k+1}$ each UAV shares the computed $[s^{k}_{i1},\dots,s^{k}_{iK}]$ (solution to Problem 3) with the UAVs that are from its neighborhood $\mathcal{N}_{ik+1}$ at time instance $t_{k+1}$. For the $j^{\rm th}$ UAV from $\mathcal{N}_{ik+1}$, we have $\bar{s}^{k+1}_{j\tau}=s^{k}_{j\tau+1}$, $\tau=1,\dots,K-1$. The algorithm starts at time $t_{1}$ ($k=1$) with $s_{i0}^{1}=\gamma_{i}^{0}$, $l_{i0}^{1}=\dot{\gamma}_{i}^{0}$ and $\bar{s}_{j\tau}^{1}=\gamma_{j}^{0}+\tau h$, $\tau=1,\dots,K$, which are known due to initial information sharing between neighboring UAVs. Eventually, as approximations to $\gamma_{i}(t)$ and to its first-order and second-order derivatives, we set
$\begin{split}\gamma_{i}(t_{k})\approx s^{k}_{i1},\quad\dot{\gamma}_{i}(t_{k})\approx\ell^{k}_{i1},\quad\ddot{\gamma}_{k}(t_{k})\approx u^{k}_{i1}.\end{split}$
The second term of the cost function (20) shows that the change in the communication can significantly impact the virtual time, particularly its second-order derivative. To mitigate and smoothen these effects, mainly in scenarios where communication depends on the distance between UAVs, we introduce the smoothing function:
$\phi\left(z,p_{1},p_{2}\right)=\begin{cases}1,&z<p_{1},\\ \eta(z),&p_{1}\leqslant z\leqslant p_{2},\\ 0,&p_{2}<z,\end{cases}$ (24)
where $p_{1}<p_{2}$ are positive parameters and $0\leqslant\eta(z)\leqslant 1$, such that $\phi\in C^{2}({\mathbb{R}}^{+})$ for any $p_{1}<p_{2}$.
The function $\phi$ quantifies the quality of communication between UAVs based on their relative distances. Communication is excellent if the distance between two UAVs is less than $p_{1}$. In contrast, communication is lost if the distance exceeds $p_{2}$. For distances in the intermediate range $\left[p_{1},p_{2}\right]$, the quality of communication smoothly transitions between these two extremes. The modified cost function is
$J_{i}\left(y^{k}_{i},\bar{s}^{k}_{-i},\phi\right)=\sum_{\tau=1}^{K}\left(\ell^{k}_{i\tau}-1\right)^{2}\\ +F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi\right)+\sum_{\tau=0}^{K-1}{u^{k}_{i\tau}}^{2}$ (25)
with
$\begin{split}F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi\right)=\sum_{j\neq i}&\phi\left(d_{i,j}^{k,0},a,b\right)\sum_{\tau=1}^{K}\left(s^{k}_{i\tau}-\bar{s}^{k}_{j\tau}\right)^{2},\end{split}$
where
$d_{i,j}^{k,\tau}=\left|\left|x_{d,i}\left(s^{k}_{i\tau}\right)-x_{d,j}\left(\bar{s}^{k}_{j\tau}\right)\right|\right|.$
Although the smoothing function is not convex, the modified cost function is quadratic because the function $\phi$ is treated as a constant in the cost function for each MPC step. Since it also does not impact the constraints, Problem 3 remains a quadratic optimization problem.
Remark 2. The formulation of Problem 3 is agile and can be modified to achieve additional goals; for example, the cost function can be augmented with additional terms that ensure collision avoidance:
$J_{i}\left(y^{k}_{i},\bar{s}^{k}_{-i},\phi,\psi\right)=\sum_{\tau=1}^{K}\left(\ell^{k}_{i\tau}-1\right)^{2}+\sum_{\tau=0}^{K-1}{u^{k}_{i\tau}}^{2}\\ +F_{i}^{ca}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi\right)+F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi,\psi\right).$ (26)
The third term of (26) is defined as follows:
$\begin{split}F_{i}^{ca}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi\right)=&\sum_{j\neq i}\sum_{\tau=1}^{K}C_{i}\frac{\phi\left(d_{i,j}^{k,\tau},a,b\right)}{\left({d_{i,j}^{k,\tau}}\right)^{2}}.\end{split}$ (27)
This term becomes active when the distance between the UAVs falls below $b$, prompting the UAVs to adjust the speed and ensure collision avoidance. The effect of the collision avoidance term (27) is the highest when the distance is less than or equal to $a$. $C_{i}$ is a positive parameter that weighs the collision avoidance term impact for each UAV. Additionally, to ensure time coordination, the term $F_{i}^{cm}$ is defined as follows:
$\begin{split}F_{i}^{cm}\left(s^{k}_{i},\bar{s}^{k}_{-i},\phi,\psi\right)=\sum_{j\neq i}\phi\left(d_{i,j}^{k,0},c,d\right)\psi\left(d_{i,j}^{k,0},a,b\right)\\ \times\sum_{\tau=1}^{K}\left(s^{k}_{i\tau}-\bar{s}^{k}_{j\tau}\right)^{2},\end{split}$
where $a<b<c<d$ and
$\psi(z,q_{1},q_{2})=\begin{cases}0,&z<q_{1},\\ \xi(z),&q_{1}\leqslant z\leqslant q_{2},\\ 1,&q_{2}\leqslant z,\end{cases}$ (28)
$q_{1}<q_{2}$ are positive parameters and $0\leqslant\xi(z)\leqslant 1$, such that $\psi\in C^{2}({\mathbb{R}}^{+})$. As a result, time coordination is achieved when the distance between the UAVs is greater than $a$. Furthermore, coordination is disregarded when the distance is less than $a$, and only collision avoidance is active.
Although the cost function (26) is not convex, the optimization problem is still simple enough to solve it efficiently in real-time.
IV Simulations
In this section we demonstrate the effectiveness and applicability of the proposed method in complex, realistic scenarios.
IV-A Simulation Framework
We use a high-fidelity simulation framework with multirotor UAV dynamics to demonstrate the performance of the algorithm. The simulation setup incorporates factors such as aerodynamic effects, actuator limitations, and environmental disturbances. The simulations validate the methodology and showcase its agility when deployed in various scenarios.
Software Setup
The simulations were performed using RotorPy [12], an open-source Python-based simulator designed for multirotor UAVs. RotorPy provides dynamic modeling of multirotor systems, including aerodynamic effects and nonlinear equations of motion.
The simulation scenarios were designed to explore different aspects of the proposed coordination algorithm. The details of the parameters concerning the MPC used in the simulation scenarios described in this section can be found in Table I. In all scenarios, we use $C_{i}=i+1$ in the collision avoidance term (27). Varying parameters for the scenarios are the communication terms $c$, $d$, and the collision avoidance terms $a$ and $b$, whose values are accordingly presented.
TABLE I: Parameters for MPC used in simulations
Parameters Value
Number of UAVs $N=6$
Prediction horizon $K=10.0$
Time step $h=0.05$ (s)
State constraint $\gamma^{i}_{min}=0.0$
State constraint $\dot{\gamma}^{i}\in\left[0.0,2.0\right]$
Control constraint $\ddot{\gamma}^{i}\in\left[-6.0,6.0\right]$
The algorithm, Alg. 1, employed in these simulations, is designed as follows: the desired state for each UAV is computed, and then passed to the path-following controller that guides the UAV. Subsequently, the path-following error is evaluated, and the initial conditions of the Problem 3 are updated. Then, based on the solution of Problem 3 the UAV transmits $[s^{k}_{i2},\dots,s^{k}_{iK}]$ to the UAVs from the neighbourhood $\mathcal{N}_{ik+1}$. Afterwards, the desired state for the UAV is updated, and the process is repeated iteratively. This process results in the coordination of the multi-agent system.
1
2Initialize: number of agents $N$, desired trajectories $x_{di}$, state and control constraints $\dot{\gamma}^{i}_{\min}$, $\dot{\gamma}^{i}_{\max}$, $\ddot{\gamma}^{i}_{\max}$, initial conditions $\gamma_{i}^{0}$, $\dot{\gamma_{i}}^{0}$, $i=1,\dots,N$, prediction horizon $K$, total time $T$, time step $h$;
3 Set $k=1$;
4
5while $kh\leqslant T$ do
6       for $i\leftarrow 1$ to $N$ do
7             $x_{i}(kh)\leftarrow\textnormal{{GetActualPosition}}(x_{di}(s_{i0}^{k}))$;
8
9            Compute $\alpha_{i}^{k}(x_{i}(kh))$;
10
11            Update initial conditions: $s_{i0}^{k}\leftarrow s^{k-1}_{i1}-\alpha_{i}^{k}$, $l_{i0}^{k}\leftarrow l_{i1}^{k-1}$;
12
13            $(s_{i}^{k},l_{i}^{k})\leftarrow\textnormal{{SolveProblem3}}(s_{i0}^{k},l_{i0}^{k},\bar{s}_{-i}^{k})$;
14
15            Transmit $[s^{k}_{i2},\dots,s^{k}_{iK}]$ with UAVs from $\mathcal{N}_{ik+1}$;
16
17            $\bar{s}^{k+1}_{i}\leftarrow[s^{k}_{i2},\dots,s^{k}_{iK}]$;
18
19      $k\leftarrow k+1$;
20
Algorithm 1 Multi-agent time-critical MPC
IV-A1 Ideal Communication and Ideal Path-Following
We begin with an ideal scenario with perfect communication and no external disturbances, where six UAVs navigate along non-overlapping circular trajectories that share a common center but have different radii. The UAVs aim to minimize the cost function (25), with each starting its mission at a distinct time $\gamma_{1}^{0}=2$, $\gamma_{2}^{0}=1$, $\gamma_{3}^{0}=0$, $\gamma_{4}^{0}=3.5$, $\gamma_{5}^{0}=4$, and $\gamma_{6}^{0}=3$. Note that only the 3rd UAV starts its mission precisely on schedule, as illustrated in Fig. 3(a). The total mission lasts 22 seconds. The results demonstrate rapid synchronization among the UAVs, making $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma_{4}=\gamma_{5}=\gamma_{6}$ in around 3.2 seconds, as shown in Fig. 3(b). Moreover, the UAVs successfully adjust their velocities to converge to their desired speed profiles (see Fig. 3(c)). The control input converges to zero by the 7th second (Fig. 4(a)). The simulation results indicate that solutions of Problem 1 are exponentially stable, which validates Theorem 3.
Figure 3: Ideal communication, ideal path following: (a) A top view of the actual trajectories followed by the UAVs under ideal conditions, with red circles marking the starting points and stars indicating the final positions; (b) $\gamma_{i}$ over time; (c) $\dot{\gamma_{i}}$ over time.
IV-A2 Non-Ideal Communication and Ideal Path-Following
Building on the conditions of the ideal communication scenario, this case introduces non-ideal communication among UAVs. The parameters for the communication term are set as follows: $c=3.5$m, $d=7$m. At the beginning of the mission, UAV 5 and UAV 6 are restricted to communicating only with the fourth one. Despite intermittent communication interruptions and subsequent reconnections, the UAVs successfully achieve coordination, with the fourth UAV serving as a communication link between all UAVs.
[…]
Figure 4: (a) Ideal communication, ideal path following, $\ddot{\gamma_{i}}$ over time, (b) Non-ideal communication, ideal path following, $\gamma_{i}$ over time, (c) Non-ideal communication, ideal path following, $\ddot{\gamma_{i}}$ over time.
IV-A3 Non-Ideal Communication and Non-Ideal Path-Following
In non-ideal path following scenario an additional wind disturbance is introduced along with low communication quality. Wind disturbance affects the motions of UAV 5 and UAV 6 at the beginning of the 22-second long simulation for 10 seconds. The wind force begins with a speed of 5 meters per second and gradually diminishes to zero by the 10-second mark.
The wind-affected trajectories, depicted in Fig. 5(a), show the impact of the path-following term $\alpha_{i}^{k}$ (see (23)) introduced in the algorithm. We set the parameter $\delta=1$ in the path-following error term. This term enables the UAVs to realign with their desired trajectories, ensuring that coordination happens during the mission. However, complete synchronization is observed later compared to the previous scenario, approximately at the 12 second mark, 2 seconds after the wind stops. The control input reduces to zero after the wind effect subsides. In particular, compared to the results of the non-ideal communication and ideal path following simulation results, the drones put more effort in order to adhere to their corresponding desired paths. In the presence of wind disturbance and poor communication quality, the stabilization of the control input is notably delayed, requiring approximately 15 seconds to converge to zero, as illustrated in Fig. 5(c). This represents a significant increase compared to the 10 seconds observed in the ideal path-following scenario, Fig. 4(c).
Figure 5: Non-ideal communication, non-ideal path following: (a) A top view of the trajectories influenced by wind disturbance, with red circles marking the starting points and stars indicating the final positions; (b) $\gamma_{i}$ over time; (c) $\ddot{\gamma_{i}}$ over time.
IV-A4 Collision Avoidance
Based on Remark 2, we demonstrate coordination along with collision avoidance, a scenario involving six UAVs following Lissajous trajectories intersecting at a single central point was implemented (Fig. 6(a)). The trajectories are given by ${\displaystyle x(k)=X\sin(vk+\epsilon)}$ and ${\displaystyle y(k)=Y\sin(wk)}$, with the following parameter values: $X=0.8$, $Y=8$, $v=0.5$, $w=0.25$, $\epsilon=0$. In this setup, no time-delays are implemented; thus $\gamma_{i}^{0}=\gamma_{j}^{0}=0$. The communication terms were set as follows: $c=10$m, $d=20$m, ensuring perfect communication throughout the mission. Collision avoidance parameters were set separately for each of the six UAVs: $a=[2.35,2.5,2.7,2.8,2.9,3.0]$m and $b=[4.7,5.0,5.4,5.6,5.8,6.0]$m. The mission lasted 42 seconds, during which the UAVs completed three passes through their trajectory, successfully avoiding collisions at the intersection point and coordinating afterward. As illustrated in Fig. 6(b), the UAVs avoided collisions between seconds 8 and 13, followed by complete synchronization between seconds 16 and 18. For further clarity, the minimal distance between any two UAVs was plotted over time, Fig. 6(c), which confirms that a minimum separation of at least 0.5 meters was consistently maintained throughout the mission, making sure that no collisions occurred during the simulation.
Figure 6: Collision avoidance under non-ideal communication and ideal path following: (a) A top view of the actual trajectories followed by the UAVs, with red circles marking the starting points and stars indicating the final positions; (b) $\gamma_{i}$ over time; (c) Minimum distance maintained between any two UAVs over time.
Scalability
The method is inherently distributed and relies only on local information, making it naturally scalable. Simulations confirm practical applicability for systems with several dozen UAVs. As the number of agents increases, computation time grows slightly (see Table II), since shared data enters the optimization as constants. However, consensus time decreases, even under partial communication loss, due to increased overall connectivity.
# of UAVs Mean Time Max Time Consensus Time
10 0.006389 0.017879 7.1
20 0.008598 0.019256 5.2
30 0.009026 0.023377 4.75
TABLE II: Mean, Max MPC step calculation times and consensus achievement time in communication failure scenario.
The simulation results show the performance of the proposed coordination algorithm in various scenarios. Under ideal conditions, UAVs achieve synchronization rapidly. Introducing non-ideal communication and wind disturbances reveals the ability of the algorithm to handle real-world challenges, ensuring coordination and maintaining safety through collision avoidance mechanisms. For reference, all simulation videos can be found in the supplementary materials.
V Flight Experiments
In this section, we demonstrate the real-time applicability of the method and highlight its communication efficiency, which enables fast information exchange and processing in critical situations. Furthermore, as mentioned before, mission operational requirements can be directly incorporated into the algorithm. In the flight experiments, collision avoidance was integrated as an example.
The flight experiments involve multiple UAVs operating cooperatively in a shared airspace. The results validate the proposed algorithm’s ability to ensure coordination and collision avoidance under realistic conditions of path-following errors and communication failures.
V-1 Non-Ideal Communication and Non-Ideal Path-Following
To validate the effectiveness of the proposed method (see Section III-B), we conducted a flight experiment involving four UAVs exchanging information. Nominal trajectories for the UAVs (Fig. 7(a) dashed circles) are designed to have overlapping regions. The collision avoidance mechanics (see Remark 2) eliminates the requirement for UAV trajectories to maintain spatial separation, as defined in (3). Consequently, UAVs can navigate overlapping paths while ensuring collision-free path-following at all times.
If the distance between UAVs is less than the communication threshold distance ($p_{2}=d$ in (24)), then they exchange information (consensus parameters $s_{i}^{k}$ and desired coordinates $x_{d,i}(s_{i}^{k})$), which enables coordination and collision avoidance via corresponding penalty terms, see (26). Since coordination and collision avoidance represent inherently contradictory objectives—where coordination cannot be maintained if UAVs must simultaneously avoid each other at the coordinates where trajectories intersect—the algorithm dynamically deactivates the coordination penalty term, activates the collision avoidance penalty term when distances between UAVs are smaller than predefined collision avoidance threshold distances ($q_{1}=a$ in (28)). During a collision avoidance maneuver, UAVs are guaranteed to communicate with each other as the collision avoidance threshold distance is always less than the communication threshold distance ($a<d$). If multiple UAVs arrive at a potential collision point with coordinated states and nominal trajectories resemble symmetry, priority is assigned to UAVs with the smallest index (this is achieved by setting $C_{i}=i$ for $i=1,2,\dots,N$ in (27)).
TABLE III: Summary of parameters used for MPC problem (Alg. 1) in flight experiment.
Parameters Value
Number of UAVs $N=4$
Prediction horizon $K=20$
Time step $h=0.2$ (s)
State constraint $\gamma_{min}^{i}=0.0$
State constraint $\dot{\gamma}^{i}\in[0.0,2.0]$
Control constraint $\ddot{\gamma}^{i}\in\left[-15.0,15.0\right]$
Comm. threshold $d=1.0,\,c=0.5$ (m)
Coll. threshold $b=0.39,\,a=0.195$ (m)
V-2 Flight Setup
UAVs follow circular flight paths, the top view of which is shown in Fig. 7(a). The blue, yellow, green, and red curves represent the actual flight trajectories, while the dashed circles represent the nominal trajectories. The UAVs begin their flight on outer circles (black dashed lines) with initial uncoordinated states (black triangles, Table IV) and gradually move to the collision point (blue square) in the clockwise direction to enter the smallest circle (dashed purple circle). After completing a single lap on the smallest circle, all UAVs exit to outer circles from the same collision point (blue square) and continue the flight on outer circles till the final point (purple triangle). Initial states and trajectory parameters are summarized in Table IV. Additionally, $\dot{\gamma}_{i}(0)=1$ for $i=1,2,3,4$, the smallest circle radius, $r_{s}=0.5$ m, the smallest circle center, $\bar{x}_{s}=\left[0.0,0.0,1.0\right]$. Furthermore, the initial velocities, accelerations, angular velocities and yaws are set to zeros, for all UAVs. The circular flight paths were designed to ensure all UAVs have the same angular rates with respect to their circle centers (Table IV), guaranteeing a collision at the collision point and during the smallest circle traversal in the absence of a collision avoidance mechanism. Lastly, we employed the smoothing function (24) within the communication term to attenuate high-frequency effects associated with the second-order derivatives of consensus parameters. This mitigation is essential to prevent potential destabilizing effects on flight performance.
TABLE IV: Initial conditions and trajectory parameters.
Initial Coordination States Initial Positions (m) Outer Circle Radius (m) Outer Circle Center (m)
$\gamma_{1}(0)=0$ $x_{1}(\gamma_{1}(0))=\left[-0.19,-0.7,1.0\right]$ $r_{1}=0.7$ $\bar{x}_{1}=\left[-0.2,0.0,1.0\right]$
$\gamma_{2}(0)=3.33$ $x_{2}(\gamma_{2}(0))=\left[-0.84,-0.77,1.0\right]$ $r_{2}=0.9$ $\bar{x}_{2}=\left[-0.4,0.0,1.0\right]$
$\gamma_{3}(0)=6.66$ $x_{3}(\gamma_{3}(0))=\left[-1.55,-0.55,1.0\right]$ $r_{3}=1.1$ $\bar{x}_{3}=\left[-0.6,0.0,1.0\right]$
$\gamma_{4}(0)=10.0$ $x_{4}(\gamma_{4}(0))=\left[-2.1,0.0,1.0\right]$ $r_{4}=1.3$ $\bar{x}_{4}=\left[-0.8,0.0,1.0\right]$
V-A Experimental Hardware and Software Setup
In this subsection, we review the hardware and software setup, while the final subsection will summarize the experimental results.
The experiment is conducted using four Crazyflie 2.1 mini quadrotors within a 5$\times$7$\times$3m Vicon motion capture arena, utilizing the Crazyswarm software framework [29]. The computational processing for the proposed algorithm is performed on a Ground Station Computer (GSC) equipped with an Intel i7-7700HQ CPU, while the Crazyflies’ integrated PID controllers handle lower-level command execution. The MPC problem, Problem 3, is solved using the CasADi IPOPT numerical solver [4] with parameters listed in Table V-1. Coordinate and orientation data are processed by the Vicon Ground Station Computer (VGSC) and streamed to GSC at 30 Hz via a wired connection. Upon receiving coordinate and orientation information from VGSC, GSC transmits it to the UAVs, along with the state command that UAVs need to follow at 10 Hz, using four Crazyradio 2.0 ISM band radio antennas.
Figure 7: (a) Top view of the flight experiment trajectories. (b) Coordination variable $\gamma_{i}(t)$ for $i=1,2,3,4$, showcasing the temporal evolution of coordination among the UAVs.
V-B Flight Experiments
The results of the flight experiment can be divided into five distinct phases. The first phase spans from $t=0$ to $t=8.6$, during which the UAVs achieve coordination starting from uncoordinated states. This is evident in Fig. 7(b), where the coordination parameters, $\gamma_{i}$, have almost equal values. Fig. 8(b) illustrates how the third and fourth UAVs decelerate ($\ddot{\gamma}_{3},\,\ddot{\gamma}_{4}<0$), while the first and the second UAVs accelerate ($\ddot{\gamma}_{1},\,\ddot{\gamma}_{2}>0$) to catch up in this time interval. Ultimately, the paces of all UAVs converge close to the nominal value ($\dot{\gamma}_{i}=1$) by the end of this phase. The second phase, from $t=8.6$ to $t=19.89$, involves the UAVs performing collision avoidance while transitioning from the outer circles to the smallest circle. This behavior is observable in Fig. 8(b), where the UAVs accelerate at varying rates to avoid collisions. Coordination is temporarily disrupted ($\gamma_{1}\neq\gamma_{2}\neq\gamma_{3}\neq\gamma_{4}$) during this phase, as reflected in Fig. 7(b). The third phase spans $t=19.89$ to $t=37.2$, during which the UAVs continue collision avoidance while gradually reestablishing coordination. By the end of this phase, the pace of the UAVs converges to the nominal values as seen in Fig. 8(a), and the coordination variables stabilize at equal offsets, ensuring collision-free path following. The fourth phase occurs between $t=37.2$ and $t=47.6$, where UAVs avoid collisions while exiting the smallest circle to return to the outer circles. Similar to phase two, the UAVs increase their speeds Fig. 8(a) by accelerating Fig. 8(b) to prevent collisions. In the final fifth phase, spanning $t=47.6$ to $t=58.6$, the UAVs regain coordination on the outer circles. It can be observed in Fig. 7(b), where coordination parameters converge to the same value ($\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma_{4}$). UAVs complete the mission earlier at $t=58.6$ than the anticipated $t=72.0$ due to maintaining higher-than-nominal pace during collision avoidance maneuvers throughout the flight.
In summary, the outcomes of the flight experiment demonstrate the effectiveness of the proposed algorithm in achieving coordination and collision avoidance among UAVs executing complex time-critical mission in the presence of communication failures. Video footage of this experiment, along with the scenario shown in Fig. 1, is available in the supplementary materials.
TABLE V: Summary of communication between UAVs during the flight experiment.
Time (s) / UAVi,j $z_{1,2}$ $z_{1,3}$ $z_{1,4}$ $z_{2,3}$ $z_{2,4}$ $z_{3,4}$
$2.290<t<t_{1}=$ 2.29
$t_{1}<t<t_{2}=2.89$
$t_{2}<t<t_{3}=14.49$
$t_{3}<t<t_{4}=52.29$
$t_{4}<t<t_{5}=58.69$
Figure 8: (a) $\dot{\gamma}_{i}(t)$ for $i=1,2,3,4$, showcasing the rate of change of the coordination variable over time. (b) $\ddot{\gamma}_{i}(t)$ for $i=1,2,3,4$, showcasing the acceleration of the coordination variable over time.
V-B1 Non-Ideal Communication and Non-Ideal Path-Following Under Wind Disturbance
We also conducted experiment with four UAVs (see Fig. 9) under wind disturbances and communication failures to evaluate the robustness of the proposed algorithm. The fourth UAV experienced the strongest headwind impact. The mean wind speed was 1.34 m/s. The UAV’s ideal speed for coordination in this scenario is 0.37 m/s, while its velocity constraint is 1.4 m/s. The coordination algorithm thus automatically adjusts the speed to enable coordination with minimum deviation.
Figure 9: Flight experiment under wind disturbance with four UAVs.
VI Limitations
In this section, we outline two main limitations of the proposed method. The theoretical analysis of the game-theoretic approach has only been conducted under ideal conditions. Althought the the simulation and experiment results demonstrate the applicability in the case of time-varying networks and the presence of path-following errors; further theoretical developments are necessary to support these results.
In the proposed approach, the desired trajectories are pre- defined and fixed, which restricts the method’s flexibility and applicability. Hence, there is a need to go beyond this assumption and consider trajectory re-planning during mission execution.
VII Conclusion
In this paper, we proposed a novel game-theoretic framework for time-critical cooperative missions of UAV systems operating over a time-varying network. Our approach introduces a distributed time-coordination mechanism that ensures agile, system-wide synchronization in dynamic and uncertain environments. The low dimensionality of the proposed optimization problem allows for effective real-time implementation while incorporating UAV dynamical constraints and mission specifications.
In the simplified scenario, the existence of a Nash equilibrium (system-wide synchronization) and exponential stability of it were proven. Furthermore, to handle more realistic scenarios MPC based algorithm was developed. Extensive simulations demonstrated the method’s effectiveness and reliability in realistic scenarios. Experiments on small UAVs validated the applicability and agility of the method in challenging conditions.
This work represents a reliable and efficient framework for UAV swarm operations, with potential applications in surveillance, disaster management, to name a few.
Acknowledgments
The authors thank Professors António Pascoal and Isaac Kaminer for their insightful discussions and valuable advice during this research. The authors also extend their thanks to Astghik Hakobyan and Andy Younes for their help with the simulation work.
The work of AM, LP and TB was supported by the Higher Education and Science Committee of MESCS RA, Research Project, 24IRF-1A001.